Problem 21.12

A silver wire is 5.2 m long and 0.52 mm in diameter.

Part A

What is its resistance?

Express your answer using two significant figures.

R =   Ω  
Answer #1

resistivity of silver=1.59*10^(-8) ohm.m

resistance=resistivity*length/area

area=pi*(diameter/2)^2=2.1237*10^(-7) m^2


then resistance=1.59*10^(-8)*5.2/(2.1237*10^(-7))=0.38932 ohms